find the number of sides of a polygon with diagonals equal to forty-five less than ten times the number of sides
Answer by stanbon(75887)

find the number of sides of a polygon with diagonals equal to forty-five less than ten times the number of sides
-------------
Let the number of sides be "n"
Then number of diagonals is nC2-n
= [n(n-1)/2]-n
= [n^2-n-2n]/2
= (n^2-3n)/2
----------------------------
Equation:
[n^2-3n]/2 = 10n-45
n^2-3n = 20n-90
n^2-23n+90 = 0
(n-5)(n-18) = 0
---
number of sides = 5 or number of sides = 18
=================
Cheers,
Stan H.